In a rectangular parallelepiped, the length and the width of the base are 12 in. and 9

in., respectively. Find the volume if the length of the diagonal of the solid is 25 inches.


1
Expert's answer
2022-06-10T10:47:21-0400

Given "a=12\ in, b=9 \ in, \sqrt{a^2+b^2+c^2}=25\ in."

Substitute


"12^2+9^2+c^2=25^2"

"c^2=400, c>0"

"c=20\ in"

"V=abc=12(9)(20)=2160({in}^3)"

2160 cubic inches